Fractions 15/39 and 45/117 are equivalent fractions. 

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

The given statement is True.
Consider the two given fraction, 15/39 and 45/117

So,
(15/39) = (45/117)
By cross multiplication,we get (15 × 117) = (45 × 39)
1,755 = 1,755



Thanks